OpenStudy (anonymous):

One more problem: A train leaves a station and travels at the rate of 40 mi/hr. Two hours later a second train leaves the station and travels at the rate of 60 mi/hr. Where will the second train overtake the first?

OpenStudy (anonymous):

from the first train (A) we have: S=So+Vo(t) S=0+40(t) after two hours (when second train leaves) train A has traveled: S=40(2) S=80 In relation to the second train (B), train A starts at 80 so, rearranging Train A's eqt: Sa = 80 + 40(t) And from train B we have: Sb = 0 + 60(t) ! The second train will overtake the first whenever Sa=Sb so, 80 + 40(t) = 60 (t) 20 (t) = 80 t = 80/20 = 4hr Ans. Is tha right?

OpenStudy (anonymous):

No, they want to know the place. Where will the second train overtake the first, not the duration.

OpenStudy (anonymous):

oh misread the question, sorry. you still have to make those calculations. With how much time they take to meet you plug it back in either equation A or B and it should give you the same answer: Sa = 80 + 40 (4) Sa = 80 + 160 = 240 mi Sb = 60 (4) = 240 mi So, The second train will overtake the first 240mi from the station